If a1 is the first of n geometric means between a and b, show that a1^n+1=a^n b
1
Expert's answer
2020-07-03T16:15:54-0400

Inserting  "n" geometric means between "a" and "b" we get the following GP series of total "(n+2)" terms


"a,a_1,a_2,...a_n,b"

Let "r" be the common ratio of this GP then "b" becomes the "(n+2)th" term of the series. So we have


"a_1=ar,..., b=ar^{n+1}"

Then


"a_1^{n+1}=(ar)^{n+1}=a^n ar^{n+1}=a^nb"

Therefore


"a_1^{n+1}=a^nb"